# Market Sensitivities and Growth Differentials Across Australian Housing Markets

### v2

@@ 3.5 Regional amplification and long-run growth differentials
We will examine price growth, where a typical house in region $r$ valued at some price $p$ at some time $t_{1}$ appreciates to $f_{r}p$ at some later time $t_{2}$ (following the price index of that region), with growth $f_{r}=e^{\mu({t_{2}})-\mu({t_{1}})}$, and compare it to growth in the national index over the same period, $f_{M}=e^{U({t_{2}})-U({t_{1}})}$. Of particular interest is the market-only approximation $\mu_{r}\approx\beta_{r}U$. The market loadings $\beta_{r}$ in Table 6 reveal substantial and persistent growth differentials across Australian housing markets. A region with $\beta_{r}>1$ amplifies national market movements, while $\beta_{r}<1$ dampens them. Over the 30-year sample period, the national market index exhibits a doubling time of approximately 10 years, corresponding to a nominal compound annual growth rate (CAGR) of 7.2%, consistent with one Reserve Bank estimates of roughly 7.25% per annum over the three decades to 2015 (kohler_vandermerwe_2015). The regional $\beta_{r}$ values transform this national trend into markedly different long-run trajectories.
@@
We do not assume that the strong growth in the market factor path itself will persist going forward into the coming decades. On internationally used price-to-income metrics, Australia’s major cities (especially Sydney) sit among the least affordable in the world (e.g. demographia2025; anzcorelogic2024), which plausibly points to slower national growth ahead (i.e., longer doubling times) than in recent decades. This future national development does not impact our conclusions, as our stability claim is narrower: the loading coefficients (e.g., $\beta_{r}$), and the way Mining and Lifestyle enter as stationary spreads around the market path, appear stable. Thus, when we map a hypothetical national change $f_{M}$ into $f_{M}^{\beta_{r}}$ for region $r$, we intend a relative scaling rule conditional on whatever national path one assumes. In contrast, naturally, the doubling times do depend on national market behavior, and apply only to the already observed data. Similarly, the overall magnitude of following uncertainty estimations also depends on the national market behavior, as the uncertainty band widths (Eq. 4) depend on the doubling time of $U$: a rapidly changing market will have narrower uncertainty bands for the bounded factors.
@@
Adding the idiosyncratic component at the 10 year horizon (column x95 total) widens the bands slightly. Brisbane (1.14), Adelaide (1.10), and ACT (1.10) exhibit the tightest total bands among major cities, while Melbourne (1.25), Sydney (1.27), and Hobart (1.30) show moderate widening. Perth (1.39) and Darwin (1.39) are outliers with substantially wider uncertainty, driven by Mining exposure in Perth’s case and idiosyncratic volatility in Darwin’s. These uncertainties widen the forecast funnel but do not alter the long-run ranking by $\beta_{r}$; the central trajectory remains intact, as is made intuitive from Figures 6, 7, and 5. Here, we chose the 10-year horizon for the uncertainties to match the observed 10 year historical national house price doubling time so as to be able to compare the columns in the table.  [^10] Here, ceteris paribus, the behavior of the market $U$ does control the overall uncertainty band widths (Eq. 4), e.g. a rapidly rising national market will have narrower uncertainty bands for the bounded factors as Market factor doubling is achieved more quickly, allowing less time for the remaining factors to evolve.
@@
Figure 8 visualizes the scenario under a national doubling ($f_{M}=2$), with the 95% band decomposed by source. Perth’s large Mining contribution (brown segment) dominates its uncertainty, arising from its positive loading ($\lambda_{r}=0.60$) on the Perth–Sydney spread, and its overall variability is large compared to the other cities, emphasizing the more volatile nature of Perth’s housing market. This spread captures the resource-sector cycle; when mining booms, the spread rises and Perth benefits directly through this positive exposure. Sydney, by contrast, has a large negative loading on Mining ($\lambda_{r}=-0.43$), meaning it moves counter-cyclically to the resource boom: as the spread rises (Perth outperforms), Sydney underperforms relative to the market factor. This opposing-phase behavior explains why both cities show strong Mining factor contributions to their variance in Figure 8, despite being on opposite ends of the spread. The Mining spread variance ($\sigma^{2}_{PS}$) is substantial (Table 5), so even moderate loadings translate into meaningful uncertainty. As might be expected, much of Perth’s non-market component is associated with the Mining factor Perhaps more surprisingly, this is also the case for Sydney, through its opposing price behavior. Also, Adelaide’s volatility is largely “explained” by processes other than the Mining and Lifestyle factors of our model, suggesting a more idiosyncratic character for this city.
@@
Melbourne exhibits the strongest fundamental growth ($\beta_{r}=1.22$) and a significant negative loading on the Lifestyle factor ($\gamma_{r}=-0.70$), making Lifestyle the dominant non-market source of variation. This negative exposure suggests that Melbourne tends to move counter-cyclically with respect to amenity-driven demand shifts. As also described in Section 2.2, this is consistent with longstanding interstate migration patterns from the southern states to Queensland (which loads highly on the Lifestyle factor; see below). Several demand-side factors provide a plausible explanation: relatively high housing costs, a colder and more variable climate, which for some households may be less attractive than warmer coastal regions, and the increased feasibility of remote work, all of which may encourage some households to relocate to more affordable coastal areas (yanotti2023). Melbourne’s relative underperformance in recent years also coincides with a period of extended COVID-19 restrictions and subsequent changes to property taxation, both of which may have influenced household location and tenure decisions.
@@
Brisbane combines high fundamental growth ($\beta_{r}=1.10$) with tight uncertainty bands (x95 total = 1.14). Its positive Lifestyle loading ($\gamma_{r}=0.39$) is consistent with Brisbane’s position as a lifestyle-migration destination, while Mining plays almost no role ($\lambda_{r}=0.06$). The regional areas, Rest of QLD ($\gamma_{r}=0.47$), Rest of NSW ($\gamma_{r}=0.38$), and Rest of Tas ($\gamma_{r}=0.54$), are dominated by positive Lifestyle loadings, as expected for amenity-rich coastal and hinterland markets, e.g. Bateman’s Bay in NSW and Sunshine Coast and Surfers Paradise in QLD. They all show strong positive exposures, while their Mining loadings remain modest, except in Rest of QLD, where mining is also a significant factor alongside lifestyle, as is expected from the large mining operations in Central Queensland, for instance the coal exploration of the Bowen Basin. The idiosyncratic component is relatively small for Rest of QLD and Rest of NSW, indicating that most of the cyclical variability there is captured by the Mining and Lifestyle factors.
@@
On current trends, the stable $\beta_{r}$ hierarchy implies widening long-run level differences: if the national factor continues to rise, higher-$\beta_{r}$ cities will outpace lower-$\beta_{r}$ regions. The same scaling works in reverse in a downturn or stagnation: for any national path $f_{M}<1$, high-$\beta_{r}$ markets (e.g., Melbourne and, to a lesser extent, Brisbane) contract more, while lower-$\beta_{r}$ regional areas are less sensitive than the capitals and the national aggregate.
@@ 4 Conclusion
We studied long-run growth differences across Australian house markets (1995–2024, houses only) using a transparent three-factor model in levels: a Market factor that absorbs the main trend common to all indexes up to a coefficient, plus two more cyclical spreads (Mining and Lifestyle) that capture mean-reverting departures from the national path. These interpretable near-orthogonal factors proxy the first three principal components of the granular local price indexes, explaining most of the variance in that dataset. This allows ARIMAX estimation. We have shown that the market factor, i.e. the city-specific scaling of the national price index, explains significant parts of the price developments in the cities. Conversely, volatility aside from the market factor is largely explained by the Mining and Lifestyle factor for the most important cities and regions, with Adelaide and ACT among the exceptions, where departures arise largely from idiosyncratic movement specific to each city.
@@
Regional market sensitivities ($\beta_{r}$) are remarkably stable in expanding windows, yielding a persistent cross-sectional ranking: large capitals track or amplify national movements, while regional areas have a damped relationship to the national market. The Mining and Lifestyle spreads behave as stationary factors: they widen forecast funnels but do not overturn the long-run ordering implied by the market sensitivity. Scenario mapping is therefore simple and transparent: for any assumed national change by a factor $f_{M}$, the market-only regional baseline is $f_{M}^{\beta_{r}}$, with spread and idiosyncratic variances adding a band. Ten-year bands are generally moderate once scaled by loadings, with wider ranges only where exposures are large (e.g., Perth to Mining) or idiosyncratic volatility is high. Decomposing the bands clarifies whether Mining, Lifestyle, or the remainder drives regional risk around the market-only path. Perth’s high volatility outside the market factor is dominated by the Mining factor. Perhaps more surprisingly, this is also the case for Sydney, through its opposing price behavior.
@@
Beyond characterizing long-run growth differentials, the factor model provides a practical framework for assessing relative valuation across markets and time, as we can identify when regional markets are historically elevated or depressed relative to their long-run paths in the context of these large-scale geographic price movements. For example, in 2024 the Rest of NSW exhibits prices higher than its Market factor trajectory would suggest, attributable to an elevated Lifestyle factor consistent with sustained post-COVID migration to amenity-rich coastal areas. The factor loadings provide objective metrics for assessing whether current local price levels reflect long-term growth in line with the national market, or variations of a more cyclical nature and of geographic specificity. This complements traditional valuation metrics (e.g., price-to-income ratios). The decomposition does not assess whether the Market factor itself is overvalued in an absolute sense (e.g., relative to income fundamentals or long-run sustainable levels). While absolute affordability measures indicate whether national or regional housing is expensive relative to fundamentals, the factor decomposition reveals whether a market is expensive or cheap relative to its own historical relationship with the national Market factor. A region can be absolutely expensive (high price-to-income) yet cyclically undervalued (negative Mining or Lifestyle contributions) simultaneously, or vice versa. Naturally, the national market itself undergoes cycles, but these are experienced by all regions to varying degrees according to their $\beta_{r}$. Unlike the Mining and Lifestyle cycles, whose effects can partially offset across regions with opposing loadings, the national Market component is systematic and cannot be diversified away by geographic spreading.
With the cyclical factors removed, Melbourne leads the major cities with significantly higher amplification of the market factor, followed by Brisbane. Sydney, Adelaide and Perth are more on-par with the national pace (upon removal of the cyclical factors), whereas the regions lag. As further growth over the coming decade is not guaranteed, this pattern could also work in reverse, where a possible national downturn would be amplified in Melbourne, with the same mechanisms that support this city’s higher $\beta$ in expansions also making it more responsive in contractions. Conversely, the regional areas (away from the cities) could be impacted less by a strong national price decline than the major cities.
@@
Given that the eight capital cities collectively account for roughly two-thirds of national house sales in our sample, the national repeat-sales index is significantly influenced by the paths of Sydney, Melbourne and, to a lesser extent, Brisbane (Table 1). International evidence suggests that housing markets in large cities are characterized by limited supply (with limited immediate scope for infill developments, particularly for houses), building restrictions, slow supply response times and geographic constraints. The resulting low elasticity allows steep house price increases driven by high demand from the pull to big cities generated by employment and education opportunities and social, cultural and recreational amenities. Migration and decreasing household size add to local population pressures (kohler_vandermerwe_2015). With housing stock and wealth concentrated in these cities, price movements tend to ripple out to surrounding regions and help lead the national cycle (e.g. Nijskens2019HotProperty; otto2007). The well-documented increasing popularity and rising house prices of big cities throughout the world (Nijskens2019HotProperty) is reflected in the higher betas of the Australian cities found in our study. Complementary to this, the sustained increases in the national price index of the last 30 years, with the most recent growth period starting around 2012 reflect the big cities combined. This arises from a long-term urbanization process, where work opportunities are increasingly concentrated in the larger cities. The regions lag behind with an increasing price gap, as reflected in their low betas.
@@
Acknowledgments: I thank Yiran Yao (Monash Master of Business Analytics) for research assistance on data processing and figure preparation.
@@
The SA2 indexes serve two purposes. First, we use them for exploratory PCA (Section 2.2) to identify interpretable factor proxies replicating sijp_francke_ree2024. Second, we aggregate them to construct the SA4-level and major city-level indexes used in our factor model estimation (Sections 3.13). Aggregation weights are proportional to number of houses involved in sales during January 2015–January 2020 (a sub-period of our data window), ensuring that larger markets contribute appropriately to regional composites.
@@
For AR(2), the characteristic polynomial is $\Phi(z)=1-\phi_{1}z-\phi_{2}z^{2}$. Stationarity holds when the roots of $\Phi(z)=0$ lie outside the unit circle (ensured in estimation). With $\phi_{2}<0$, near-boundary values ($\phi_{1}\approx 2\sqrt{-\phi_{2}}$) correspond to highly persistent, gently damped cycles. The MA(1) term $\theta_{1}<0$ absorbs short-run serial correlation, and the intercept is near zero because the factor series are mean-centered by construction. Note that $\sigma$ (innov) pertains to the one-step innovation in the state-space representation and is distinct from the 10-year forecast dispersion used in the scenario funnels.

### v3

@@ 3.5 Regional amplification and long-run growth differentials
We now examine price growth, where a typical house in region $r$ valued at some price $p$ at some time $t_{1}$ appreciates to $f_{r}p$ at some later time $t_{2}$ (following the price index of that region), with growth $f_{r}=e^{\mu({t_{2}})-\mu({t_{1}})}$, and compare it to growth in the national index over the same period, $f_{M}=e^{U({t_{2}})-U({t_{1}})}$. Of particular interest is the market-only approximation $\mu_{r}\approx\beta_{r}U$. The market loadings $\beta_{r}$ in Table 6 reveal substantial and persistent growth differentials across Australian housing markets. A region with $\beta_{r}>1$ amplifies national market movements, while $\beta_{r}<1$ dampens them. Over the 30-year sample period, the national market index exhibits a doubling time of approximately 10 years, corresponding to a nominal compound annual growth rate (CAGR) of 7.2%, consistent with one Reserve Bank estimates of roughly 7.25% per annum over the three decades to 2015 (kohler_vandermerwe_2015). The regional $\beta_{r}$ values transform this national trend into markedly different long-run trajectories.
@@
Naturally, it remains unknown whether the strong growth in the market factor path itself will persist into the coming decades. On internationally used price-to-income metrics, Australia’s major cities (especially Sydney) sit among the least affordable in the world (e.g. demographia2025; anzcorelogic2024), which plausibly points to slower national growth ahead (i.e., longer doubling times) than in recent decades. This future national development does not impact our conclusions, as our stability claim is narrower: the loading coefficients (e.g., $\beta_{r}$), and the way Mining and Lifestyle enter as stationary spreads around the market path, appear stable. Thus, when we map a hypothetical national change $f_{M}$ into $f_{M}^{\beta_{r}}$ for region $r$, we intend a relative scaling rule conditional on whatever national path one assumes. In contrast to the loadings, the uncertainty band widths related to the Mining, Lifestyle and idiosyncratic factors (Eq. 4, and discussed below) associated with a national doubling do depend on the Market factor as they depend on the doubling time: a rapidly changing market will have narrower bands for the bounded factors (as doubling is achieved sooner, so that spread uncertainty has had less time to build).
@@
Adding the idiosyncratic component at the 10 year horizon (column x95 total) widens the bands slightly. Brisbane (1.14), Adelaide (1.10), and ACT (1.10) exhibit the tightest total bands among major cities, while Melbourne (1.25), Sydney (1.27), and Hobart (1.30) show moderate widening. Perth (1.39) and Darwin (1.39) are outliers with substantially wider uncertainty, driven by Mining exposure in Perth’s case and idiosyncratic volatility in Darwin’s. Here, we chose the 10-year horizon for the uncertainties to match the observed 10 year historical national house price doubling time so as to be able to compare the columns in the table.  [^11] Here, ceteris paribus, the behavior of the market $U$ does control the overall uncertainty band widths (Eq. 4), e.g. a rapidly rising national market will have narrower uncertainty bands for the bounded factors as Market factor doubling is achieved more quickly, allowing less time for the remaining factors to evolve.
@@
Figure 8 visualizes the scenario under a national doubling ($f_{M}=2$), with the 95% band decomposed by source (based on a 10 year horizon corresponding to the 30-year historical doubling time). Perth’s large Mining contribution (brown segment) dominates its uncertainty, arising from its positive loading ($\lambda_{r}=0.60$) on the Perth–Sydney spread, and its overall variability is large compared to the other cities, emphasizing the more volatile nature of Perth’s housing market. This spread captures the resource-sector cycle; when mining booms, the spread rises and Perth benefits directly through this positive exposure. Sydney, by contrast, has a large negative loading on Mining ($\lambda_{r}=-0.43$), meaning it moves counter-cyclically to the resource boom: as the spread rises (Perth outperforms), Sydney underperforms relative to the market factor. This opposing-phase behavior explains why both cities show strong Mining factor contributions to their variance in Figure 8, despite being on opposite ends of the spread. Melbourne exhibits the strongest fundamental growth ($\beta_{r}=1.22$) and a significant negative loading on the Lifestyle factor ($\gamma_{r}=-0.70$), making Lifestyle the dominant non-market source of variation. This negative exposure suggests that Melbourne tends to move counter-cyclically with respect to amenity-driven demand shifts. As also described in Section 2.2, this is consistent with longstanding interstate migration patterns from the southern states to Queensland (which loads highly on the Lifestyle factor; see below). Several demand-side factors provide a plausible explanation: relatively high housing costs, a colder and more variable climate, which for some households may be less attractive than warmer coastal regions, and the increased feasibility of remote work, all of which may encourage some households to relocate to more affordable coastal areas (yanotti2023). Melbourne’s relative underperformance in recent years also coincides with a period of extended COVID-19 restrictions and subsequent changes to property taxation, both of which may have influenced household location and tenure decisions.
@@
Brisbane combines high fundamental growth ($\beta_{r}=1.10$) with tight uncertainty bands (x95 total = 1.14). Its positive Lifestyle loading ($\gamma_{r}=0.39$) puts this city with a warmer climate in contrast to Melbourne, and is consistent with its position as a lifestyle-migration destination. Mining plays almost no role ($\lambda_{r}=0.06$). Important other recipients of the lifestyle migration push out of the southern cities are the regional areas Rest of QLD ($\gamma_{r}=0.47$), Rest of NSW ($\gamma_{r}=0.38$), and Rest of Tas ($\gamma_{r}=0.54$). They are dominated by positive Lifestyle loadings, as expected for amenity-rich coastal and hinterland markets, e.g. Bateman’s Bay in NSW and Sunshine Coast and Surfers Paradise in QLD. They all show strong positive exposures, while their Mining loadings remain modest, except in Rest of QLD, where mining is also a significant factor alongside lifestyle, as is expected from the large mining operations for instance in Central Queensland. The idiosyncratic component is relatively small for Rest of QLD and Rest of NSW, indicating that most of the cyclical variability there is captured by the Mining and Lifestyle factors. In contrast to most of the cities, Adelaide’s volatility largely arises processes other than the Mining and Lifestyle factors of our model, suggesting a more idiosyncratic character for this city.
@@
On current trends, the stable $\beta_{r}$ hierarchy implies widening long-run level differences: if the national factor continues to rise, higher-$\beta_{r}$ cities will outpace lower-$\beta_{r}$ regions. The same scaling works in reverse in a downturn or stagnation: for any national path $f_{M}<1$, high-$\beta_{r}$ markets (e.g., Melbourne) contract more, while lower-$\beta_{r}$ regional areas are less sensitive than the capitals and the national aggregate.
@@ 4 Conclusion
We studied long-run growth differences across Australian house markets (1995–2024, houses only) using a transparent three-factor model in levels: a Market factor that absorbs the main trend common to all indexes up to a coefficient, plus two more cyclical spreads (Mining and Lifestyle) that capture mean-reverting departures from the national path. These interpretable near-orthogonal factors proxy the first three principal components of the granular local price indexes, explaining most of the variance in that dataset. This near-independence allows straightforward ARIMAX estimation and interpretation. We have shown that the market factor, i.e. the city-specific scaling of the national price index, explains significant parts of the price developments in the cities. Conversely, volatility beyond the market factor is largely explained by the Mining and Lifestyle factor for the most important cities and regions, with Adelaide and ACT among the exceptions, where departures arise largely from idiosyncratic movement specific to each city.
@@
Regional market sensitivities ($\beta_{r}$) are stable in expanding windows, yielding a persistent cross-sectional ranking: large capitals track or amplify national movements, while regional areas have a damped relationship to the national market. The Mining and Lifestyle spreads behave as stationary factors: they widen forecast fans but do not overturn the long-run ordering implied by the market sensitivity. Scenario mapping is therefore simple and transparent: for any assumed national change by a factor $f_{M}$, the market-only regional baseline is $f_{M}^{\beta_{r}}$, with spread and idiosyncratic variances adding a band. Ten-year bands are generally moderate once scaled by loadings, with wider ranges only where exposures are large (e.g., Perth to Mining) or idiosyncratic volatility is high. Decomposing the bands clarifies whether Mining, Lifestyle, or the remainder drives regional risk around the market-only path. Specifically, Perth shows high volatility outside the market factor, dominated by the Mining factor. Perhaps more surprisingly, this is also the case for Sydney, through its opposing price behavior.
@@
Beyond characterizing long-run growth differentials, the factor model provides a practical framework for assessing relative valuation across markets and time, as we can identify when regional markets are historically elevated or depressed relative to their long-run paths in the context of these large-scale geographic price movements. For example, in 2024, the Rest of NSW exhibits prices higher than its Market factor trajectory would suggest, attributable to an elevated Lifestyle factor consistent with sustained post-COVID migration to amenity-rich coastal areas. The factor loadings provide objective metrics for assessing whether current local price levels reflect long-term growth in line with the national market, or variations of a more cyclical nature and of geographic specificity. This complements traditional valuation metrics (e.g., price-to-income ratios). The decomposition does not assess whether the Market factor itself is overvalued in an absolute sense (e.g., relative to income fundamentals or long-run sustainable levels). While absolute affordability measures indicate whether national or regional housing is expensive relative to fundamentals, the factor decomposition reveals whether a market is expensive or cheap relative to its own historical relationship with the national Market factor. Naturally, the national market itself undergoes cycles, but these are experienced by all regions to varying degrees according to their $\beta_{r}$. Unlike the Mining and Lifestyle cycles, whose effects can partially offset across regions with opposing loadings, the national Market component is systematic and cannot be diversified away by geographic spreading.
@@
With the cyclical factors removed, Melbourne leads the major cities with significantly higher amplification of the market factor, followed by Brisbane. Sydney, Adelaide and Perth are more on-par with the national pace (upon removal of the cyclical factors), whereas the regions lag. Given that the eight capital cities collectively account for roughly two-thirds of national house sales in our sample, the national repeat-sales index is significantly influenced by the paths of Sydney, Melbourne and, to a lesser extent, Brisbane (Table 1). International evidence suggests that housing markets in large cities are characterized by limited supply (with limited immediate scope for infill developments, particularly for houses), building restrictions, slow supply response times and geographic constraints. The resulting low elasticity allows steep house price increases driven by high demand from the pull to big cities generated by employment and education opportunities and social, cultural and recreational amenities. Migration and decreasing household size add to local population pressures (kohler_vandermerwe_2015). With housing stock and wealth concentrated in these cities, price movements tend to ripple out to surrounding regions and help lead the national cycle (e.g. Nijskens2019HotProperty; otto2007). The well-documented increasing popularity and rising house prices of big cities throughout the world (Nijskens2019HotProperty) is reflected in the higher betas of the Australian cities found in our study. Complementary to this, the sustained increases in the national price index of the last 30 years, with the most recent growth period starting around 2012 reflect the big cities combined. This arises from a long-term global urbanization process, where work opportunities are increasingly concentrated in the larger cities. The regions lag behind with an increasing price gap, as reflected in their low betas.
@@
Several limitations of our approach suggest natural extensions. We impose a low-dimensional, three-factor model with interpretable proxies for Mining and Lifestyle; while this matches the leading principal components of the granular indexes, additional factors or direct use of the PCA series could capture further geographic nuances. The analysis is based on repeat-sales house indexes, so units (apartments), may exhibit different sensitivities. Finally, relating the estimated factor loadings and time series to fundamentals such as interest rates, incomes and supply, is a natural direction for future work.
@@
Acknowledgments: I thank Yiran Yao (Monash Master of Business Analytics) for research assistance, text editing, stimulating discussions, data processing and creating Appendix D during her internship at NEOVAL. I also thank the participants of the 2025 Time Series and Forecasting Symposium at the University of Sydney for their helpful feedback on my presentation of this study.
@@
The SA2 indexes serve two purposes. First, we use them for exploratory PCA (Section 2.2) to identify interpretable factor proxies replicating sijp_francke_ree2024. Second, we aggregate them to construct the SA4-level and major city-level indexes used in our factor model estimation (Sections 3.13). Aggregation weights are proportional to number of houses involved in sales during January 2015–January 2020 (a sub-period of our data window, selected to balance stability in transaction volumes with proximity to current market conditions), ensuring that larger markets contribute appropriately to regional composites.
@@
For AR(2), the characteristic polynomial is $\Phi(z)=1-\phi_{1}z-\phi_{2}z^{2}$. Stationarity holds when the roots of $\Phi(z)=0$ lie outside the unit circle (ensured in estimation). With $\phi_{2}<0$, near-boundary values ($\phi_{1}\approx 2\sqrt{-\phi_{2}}$) correspond to highly persistent, gently damped cycles. The MA(1) term $\theta_{1}<0$ absorbs short-run serial correlation, and the intercept is near zero because the factor series are mean-centered by construction. Note that $\sigma$ (innov) pertains to the one-step innovation in the state-space representation and is distinct from the 10-year forecast dispersion used in the scenario fans.